OpenStudy (anonymous):

What is the solution to the equation 7^3x ≈ 9 ? x = 0.376 x = 0.295 x = –0.295 x = –0.376

OpenStudy (anonymous):

is it 73 or something else

OpenStudy (anonymous):

7^3x

OpenStudy (anonymous):

so what us 7^3

OpenStudy (anonymous):

use calculator

OpenStudy (anonymous):

0.376

OpenStudy (anonymous):

7^3=343

OpenStudy (anonymous):

3x ln(9)= ln(7) 3x = ln(7)/ln(9) x= 1/3 (ln(7)/ln(9))

OpenStudy (anonymous):

343/9=38.1

OpenStudy (anonymous):

x = 0.295

OpenStudy (anonymous):

now divide 9/wht u got

OpenStudy (anonymous):

so x=376

OpenStudy (anonymous):

You take log with the base seven on both sides to get \[\log_{7} \times 9=3x\]

OpenStudy (anonymous):

0.376